OpenStudy (anonymous):

A community organizes a phone tree in order to alert each family of emergencies. In the first stage, one person calls five families. In the second stage, each of the five families calls another five families, and so on. How many stages need to be reached before 19,530 families or more have been called?

OpenStudy (anonymous):

More formulas to be used. Is it recursive?

OpenStudy (anonymous):

\[a_{n} = (5)(19530)^{n - 1}\]

OpenStudy (ddcamp):

Not quite recursive, but you would have the sum: \[\sum_{n=0}^{A} 5^n = 1 + 5 + 25 + 125 + 625 +... +5^A\]

OpenStudy (anonymous):

(5n + 1) = 1 + 5 + 25 + 125 + 625? 781 stages?

OpenStudy (anonymous):

5n - 780 = 0 n = 156?

OpenStudy (ddcamp):

At the 0 stage, 1 person has been "contacted". At the 1st stage, 6 people (the 5 new and the original) have been contacted. At the second stage, the first 6 and 25 new people have been contacted. on the nth stage, you have the the people from before plus 5^n new people. Which power of 5 is closest to 19530?

OpenStudy (phi):

use the sum of a geometric series https://en.wikipedia.org/wiki/Geometric_series#Formula

OpenStudy (phi):

in your case \[\sum_{k=0}^{N}ar^k = a \frac{ 1-r^{N+1 }}{ 1-r }\] a=1 and r= 5 and you want to solve for N

OpenStudy (phi):

in other words, solve for N in \[ \frac{1 -5^{N+1}}{-4}= 19530 \]

OpenStudy (anonymous):

-3125?

OpenStudy (phi):

the first step is multiply both sides by -4 \[ -4 \cdot \frac{1 -5^{N+1}}{-4}= 19530 \cdot -4 \\ 1 -5^{N+1} = - 78120 \] add -1 to both sides \[ -1 + 1 - -5^{N+1} = - 78120 -1 \\-5^{N+1} = - 78121\] multiply both sides by -1 \[ 5^{N+1} = 78121\]

OpenStudy (phi):

to find N+1, take the log of both sides \[ \log(5^{N+1}) = \log(78121) \]

OpenStudy (phi):

use the property \[ \log(a^b) = b \log(a) \]

OpenStudy (phi):

\[ \log(5^{N+1}) = \log(78121) \\ (N+1) \log(5) = \log(78121) \] divide both sides by log 5 \[ N+1 = \frac{\log(78121)}{\log(5) } \] add -1 to both sides \[ N= \frac{\log(78121)}{\log(5) } -1 \]

OpenStudy (phi):

you will need a calculator to find N

OpenStudy (anonymous):

5.9999 n = 6? \[\frac{ 7815 }{ 2 }\]

OpenStudy (phi):

yes, n=6

OpenStudy (anonymous):

\[3907\frac{ 1 }{ 2 }\]

OpenStudy (phi):

what is 3907 ½ ?

OpenStudy (anonymous):

Mixed fraction of 7815. 6 stages need to be reached before 19,530?

OpenStudy (anonymous):

How can I double check that?

OpenStudy (anonymous):

\[\frac{1 -5^{6+1}}{-4}\]

OpenStudy (phi):

In the first stage, one person calls five families. that means 1+5 know after 1 stage In the second stage, each of the five families calls another five families, that means 1+ 5 + 25 know after 2 stages 6 stages will have 7 numbers: 1+5+25+125+625+3125+15625 that we add up

OpenStudy (phi):

If we did not use the equation, we could solve the problem by keep adding stages and adding up as we go until we pass 19530

OpenStudy (anonymous):

It's like a shortcut instead of adding all those numbers up?

OpenStudy (phi):

yes. For 7 numbers we don't really need it.. but it is a good formula to know.

OpenStudy (anonymous):

Thank you @DDCamp as well.